40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5011     Entry time: Thu Jul 21 10:08:30 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: sus sensor summary 

OSEM voltages to be corrected at upcoming vent: threshold ~ 0.7-1.2V, ( at 22 out of 50 )

ITMX_UL, UR, LL, LR, SD

ITMY_UL

ETMX_UL, UR, LL, LR, SD

ETMY_UL

SRM_UL, UR, LL

MC2_UR_SD

BS_UR, SD

MC3_UL, LR, LL

 

Attachment 1: sus_sensor.png  42 kB  | Hide | Hide all
sus_sensor.png
ELOG V3.1.3-